OpenStudy (anonymous):

If the sides of the square are 10 cm: What is the area of the square? What is the area of the circle? What is the area of the blue section?

OpenStudy (yoloshroom):

a^2 = area of a square so 10^2 = area

OpenStudy (anonymous):

Find the area of a square by multiplying 10*10. = 100cm^2 Find the area of a circle using pi r^2. The radius is half of the diameter so the radius is 5. Subtract the area you get for the circle from the area of the square to get the area in the blue section. So (Area of square) - (Area of circle) = (Area of blue section)

OpenStudy (anonymous):

so area is 78.54

OpenStudy (anonymous):

Of the circle. yes :). Now do 100 - 78.54 to find the area of the blue.

OpenStudy (anonymous):

so its 21.46

OpenStudy (anonymous):

could you help me with another one? @Loopki40